OpenStudy (anonymous):

Solve 2x2 - 8x = -7. I just want to know if my answer is correct. I got x=2+- 2 sqr 2

OpenStudy (swissgirl):

\(2x^2-8x+7=0\) \(A=2, B=-8, C=7\) \(\large \frac{-(-8) \pm \sqrt{(-8)^2-4(2)(7)}}{2(2)}\) \(\large \frac{8 \pm \sqrt{64-56}}{4}\) \(\large \frac{8 \pm \sqrt{8}}{4}\) \(\large \frac{8 \pm 2\sqrt{2}}{4}\) \(\large \frac{4\pm \sqrt{2}}{2}\) OR \(\large 2\frac{\pm \sqrt{2}}{2}\)

OpenStudy (swissgirl):

hmmm are you sure this is the right equation?

OpenStudy (anonymous):

oh sorry these are the answer choices: negative 2 plus or minus square root of 2 negative 2 plus or minus 2 square root of 2 quantity of 2 plus or minus square root of 2 all over 2 2 plus or minus square root of 2 end root over 2

OpenStudy (swissgirl):

So what do you think?

OpenStudy (anonymous):

im not sure I cant understand what you typed up there..haha

OpenStudy (swissgirl):

So how did you get your answer?

OpenStudy (anonymous):

my own answer is x=2+-2 sqr 2

OpenStudy (swissgirl):

How did you get that?

OpenStudy (anonymous):

its long in my notebook haha I just need to know if my answer is correct or not

OpenStudy (swissgirl):

Your answer is incorrect

OpenStudy (swissgirl):

Look at my answer above

OpenStudy (anonymous):

OH it shows now! thanks! ill review it and find my mistake!

OpenStudy (swissgirl):

\( \large 2 \pm \frac{\sqrt{2}}{2}\)

OpenStudy (anonymous):

Thanks!